# Character Theory for Semilinear Representations

### v2

@@ 8.4.2. Action of $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{F}(H)$
For $g\in G_{e}$ and $W\in\operatorname{\mathbf{Rep}}_{F}(H)$, $g*W\in\operatorname{\mathbf{Rep}}_{F}(H)$ is the $F[H]$-module with
@@
For any $W\in\operatorname{\mathbf{Rep}}_{F}(H)$ there is a $F[H]$-module isomorphism
@@
8.5. Description of Irreducibles
We can use Corollary 8.14 to obtain a precise description of the restriction $e\cdot V$ for irreducible $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Recall that $\operatorname{Irr}_{L}^{\rtimes}(G)$ and $\operatorname{Irr}_{F}(H)$ denote the sets of isomorphism classes of irreducible objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{F}(H)$ respectively.
@@ 8.5. Description of Irreducibles
For $W\in\operatorname{Irr}_{F}(H)$, we write $\Gamma_{e,W}\coloneqq\operatorname{Stab}_{\Gamma_{e}}(W)$.
@@
Suppose that $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and $W$ is any irreducible $F[H]$-submodule of $e\cdot V$. Then
$e\cdot V=\bigoplus_{g\in G_{e}/H}(g*W)^{m(V)}$
for some $m(V)\geq 1$. Furthermore, if $W\in\operatorname{Irr}_{F}(H)$ then
  • $W$ appears in $e\cdot V$ for a unique $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$,
@@
Sending $W$ to this unique $V$ defines a bijective correspondence
$\operatorname{Irr}_{L}^{\rtimes}(G)\xleftrightarrow{\sim}\operatorname{Irr}_{F}(H)/\Gamma_{e},$
where conversely $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$ is sent to any irreducible component of $e\cdot V$.
@@
For the first claim, suppose that $W$ is any irreducible $F[H]$-submodule of $e\cdot V$, which is semisimple by Corollary 8.14. For any $g\in G_{e}$, the number of copies of each gWg*W in $e\cdot V$ must be equal to the number of copies of $W$, because the action $\rho(g)$ of $g$ on $V$ defines an isomorphism $\rho(g)\colon g*(e\cdot V)\xrightarrow{\sim}e\cdot V$. It therefore remains to see that only one orbit of the action of $G_{e}$ on $\operatorname{Irr}_{F}(H)$ appears in $e\cdot V$. Suppose then that $U\in\operatorname{Irr}_{F}(V)$, and recall that $(L\rtimes G)\otimes_{L[H]}(e\cdot V)\cong V^{\oplus|\Gamma_{e}|}$ by Corollary 8.5, so by Lemma 2.14 both $(L\rtimes G)\otimes_{L[H]}W$ and $(L\rtimes G)\otimes_{L[H]}U$ are isomorphic to a direct sum of a number of copies of $V$. In particular, there are $a,b\geq 1$ with
$\left((L\rtimes G)\otimes_{L[H]}W\right)^{\oplus a}\cong\left((L\rtimes G)\otimes_{L[H]}U\right)^{\oplus b}$
@@
$\bigoplus_{x\in G_{e}/H}(x\cdot W)^{\oplus a}\cong\bigoplus_{y\in G_{e}/H}(y\cdot U)^{\oplus b}$
by Proposition 8.17 and therefore by Lemma 2.14 $U\cong g\cdot W$ for some $g\in G_{e}$.
For the remaining claims, given this decomposition of $e\cdot V$,
$V^{|\Gamma_{e}|}\cong(L\rtimes G)\otimes_{L[H]}(e\cdot V)\cong\bigoplus_{g\in G_{e}/H}((L\rtimes G)\otimes_{L[H]}(g*W))^{m(V)}.$
In particular, each $(L\rtimes G)\otimes_{L[H]}(g*W)=V^{\oplus r}$ where $r=|\Gamma_{e,W}|/m(V)$. In particular, $W$ appears in $e\cdot V$ for exactly one $V\in\operatorname{Irr}_{L}^{\rtimes}(V)$. ∎
In general, we will see that any positive integer can occur as some $m(V)$ (Section 16). One case when we know that the numbers $m(V)$ are equal to $1$ is the following. This includes the case when $K$ is a finite field, by Wedderburn’s Little Theorem [milneCFT, Thm. IV.4.1].
@@
and therefore $\operatorname{End}_{F[H]}(e\cdot V)$ is commutative. Therefore, as $e\cdot V$ is semisimple, $e\cdot V$ cannot contain any factors with multiplicity greater than $1$, so $m(V)=1$. ∎
@@ 8.6. The Case When $L$ is a Field
$\operatorname{Irr}_{L}^{\rtimes}(G)$ $\xleftrightarrow{\sim}\operatorname{Irr}_{L}(H)/\Gamma,$
$\operatorname{Irr}_{L}^{\rtimes}(S)$ $\xleftrightarrow{\sim}\operatorname{Irr}_{L}(H)/\Gamma_{S}.$
@@
Let $\mathcal{A}$} be an orbit of $G$ on $\operatorname{Irr}_{L}(H)$, corresponding to $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and let
@@
be a decomposition of $\mathcal{A}$} into $S$-orbits. Let $U_{i}\in\operatorname{Irr}_{L}^{\rtimes}(S)$ correspond to ${\mathcal{A}}_{i}$, so
@@
for some $m(U_{i})\geq 1$ (by Theorem 8.20). Then
$V|_{S}\cong\bigoplus_{i=1}^{r}U_{i}^{\oplus\frac{m(V)}{m(U_{i})}}$
@@
for any $X\in{\mathcal{A}}_{i}$. In particular, $m(U_{i})\mid m(V)$ and $m(V)\mid m(U_{i})\cdot[\Gamma_{X}:\Gamma_{S,X}]$.
@@
We first show that $V|_{S}$ is semisimple. By Corollary 8.5,
$(V|_{S})^{|\Gamma_{S}|}\cong\operatorname{Ind}_{H}^{S}V|_{H},$
and $V|_{H}$ is semisimple by Corollary 8.14. Now $(V|_{S})^{|\Gamma_{S}|}$ is semisimple, as $\operatorname{Ind}_{H}^{S}(-)$ preserves semisimplicity, again by Corollary 8.14. Then $V|_{S}$ is semisimple, being a submodule of a semisimple module. Now let $U$ be any irreducible $L\rtimes S$-module. By Corollary 8.7, we have that
$\dim_{F}\operatorname{Hom}_{L\rtimes S}(U,V|_{S})=\dim_{L}\operatorname{Hom}_{L[H]}(U|_{H},V|_{H}),$
which is zero for $U\not\in\{U_{1},...,U_{r}\}$, hence
$V|_{S}\cong\bigoplus_{i=1}^{r}U_{i}^{a_{i}}$
for some $a_{i}\geq 1$. Using Corollary 8.7 we can compute
$a_{i}\cdot\dim_{F}(\operatorname{End}_{L\rtimes S}(U_{i}))$ $=\dim_{F}\operatorname{Hom}_{L\rtimes S}(U_{i},V|_{S}),$
$=\dim_{L}\operatorname{Hom}_{L[H]}(U_{i}|_{H},V|_{S}),$
$=\dim_{L}\operatorname{Hom}_{L[H]}\left(\bigoplus_{X\in{\mathcal{A}}_{i}}X^{\oplus m(U_{i})},\bigoplus_{X\in{\mathcal{A}}}X^{\oplus m(V)}\right),$
$=\sum_{X\in{\mathcal{A}}_{i}}\dim_{L}(\operatorname{End}_{L[H]}(X))\cdot m(U_{i})\cdot m(V).$
Again, using Corollary 8.7, we similarly have
$\dim_{F}\operatorname{End}_{L\rtimes S}(U_{i})$ $=\dim_{L}\operatorname{End}_{L[H]}(U_{i}|_{H}),$
$=\sum_{X\in{\mathcal{A}}_{i}}\dim_{L}(\operatorname{End}_{L[H]}(X))\cdot m(U_{i})^{2},$
and therefore combining,
$a_{i}\cdot m(U_{i})=m(V),$
which gives the first statement. For the second, take any $X\in{\mathcal{A}}_{i}$, note that by Theorem 8.20
$V^{\oplus\frac{|\Gamma_{X}|}{m(V)}}$ $\cong\operatorname{Ind}_{H}^{G}(X),$
$\cong\operatorname{Ind}_{S}^{G}(\operatorname{Ind}_{H}^{S}(X)),$
$\cong\operatorname{Ind}_{S}^{G}(U_{i})^{\oplus\frac{|\Gamma_{S,X}|}{m(U_{i})}}.$
In particular, by Lemma 2.14,
$\operatorname{Ind}_{S}^{G}(U_{i})\cong V^{\oplus\frac{|\Gamma_{X}|\cdot m(U_{i})}{|\Gamma_{S,X}|\cdot m(V)}}=V^{\oplus\frac{[\Gamma_{X}:\Gamma_{S,X}]\cdot m(U_{i})}{m(V)}},$
as required. ∎
@@
The main case of interest is the following. To state it, note that if $S$ is a normal subgroup of $G$, then there is an action of $G/S$ on $\operatorname{Irr}_{L}^{\rtimes}(S)$, where for $g\in G$, and $U\in\operatorname{Irr}_{L}^{\rtimes}(S)$, gUg*U is the abelian group $U$ but with $L\rtimes S$-module structure
@@
for $\lambda\in L$, $s\in S$ and $u\in U$.
@@
The action of $G/H$ on $\operatorname{Irr}_{L}(H)$ is the restriction to the diagonal of an action of the product $G/H\times G/H$ on $\operatorname{Irr}_{L}(H)$ (where one copy acts by conjugation in $G$ and the other acts on scalars - see Section 14). When $S>H$, however, this action of $G/S$ on $\operatorname{Irr}_{L}^{\rtimes}(S)$ is not the restriction of an action of the product $G/S\times G/S$ in the same way: one only has an action of the diagonal.
@@
Suppose that $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and that for some corresponding $W\in\operatorname{Irr}_{L}(H)$, $G_{W}$ is normal in $G$. Set $G_{V}\coloneqq G_{W}$, which is independent of the choice of $W$. Let $U\in\operatorname{Irr}_{L}^{\rtimes}(G_{V})$ correspond to $W$. Then gUg*U corresponds to gWg*W,
@@
and for any $g\in G$,
$\operatorname{Ind}_{G_{V}}^{G}(g*U)$ $\cong V,$
@@
Now we are interested in the condition that $L$ is furthermore a splitting field for $W\in\operatorname{Irr}_{L}(H)$ (meaning that $\operatorname{End}_{L[H]}(W)=L$, or equivalently that $W_{L^{\prime}}$ is irreducible for any field extension $L^{\prime}/L$).
@@
Suppose that $W\in\operatorname{Irr}_{L}(H)$, corresponding to $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$. Then the division $K$-algebra $\operatorname{End}_{L\rtimes G}(V)$ is central over $K$ if and only if $\Gamma$ fixes $W$ and $Z(\operatorname{End}_{L[H]}(W))=L$.

### v3

@@ 8.4.2. Action of $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{F}(H)$
For $g\in G_{e}$ and $W\in\operatorname{\mathbf{Rep}}_{F}(H)$, $g*W\in\operatorname{\mathbf{Rep}}_{F}(H)$ is $W$ as an abelian group, with
@@
For any $W\in\operatorname{\mathbf{Rep}}_{F}(H)$ there is an isomorphism
@@
8.5. Description of Indecomposables
We can use Corollary 8.14 to obtain a precise description of the restriction $e\cdot V$ for indecomposable $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Recall that $\operatorname{Ind}_{L}^{\rtimes}(G)$ and $\operatorname{Ind}_{F}(H)$ denote the sets of isomorphism classes of indecomposable objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{F}(H)$ respectively.
@@ 8.5. Description of Indecomposables
For $W\in\operatorname{Ind}_{F}(H)$, we write $\Gamma_{e,W}\coloneqq\operatorname{Stab}_{\Gamma_{e}}(W)$.
@@
Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$, and $W$ is any indecomposable direct summand of $e\cdot V$. Then
$e\cdot V=\bigoplus_{\gamma\in\Gamma_{e}/\Gamma_{e,W}}(\gamma*W)^{m(V)}$
for some $m(V)\geq 1$. Furthermore, if $W\in\operatorname{Ind}_{F}(H)$ then
  • $W$ appears in $e\cdot V$ for a unique $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$,
  • This $V$ is irreducible if and only if $W$ is irreducible,
@@
In particular, $m(V)\mid|\Gamma_{e,W}|$, and sending $W$ to this unique $V$ defines bijective correspondences
$\operatorname{Ind}_{L}^{\rtimes}(G)\xleftrightarrow{\sim}\operatorname{Ind}_{F}(H)/\Gamma_{e},\qquad\operatorname{Irr}_{L}^{\rtimes}(G)\xleftrightarrow{\sim}\operatorname{Irr}_{F}(H)/\Gamma_{e}$
where conversely $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ is sent to any indecomposable direct summand of $e\cdot V$.
@@
For the first claim, write $e\cdot V$ as a direct sum of indecomposable objects of $\operatorname{\mathbf{Rep}}_{L}(H)$ using Lemma 2.14, and let $W$ be an indecomposable direct summand of $e\cdot V$. For any $g\in G_{e}$, the number of copies of each gWg*W in $e\cdot V$ must be equal to the number of copies of $W$, using Lemma 2.14 and the fact that the action $\rho(g)$ of $g$ on $V$ defines an isomorphism $\rho(g)\colon g*(e\cdot V)\xrightarrow{\sim}e\cdot V$.
It therefore remains remains to see that only one orbit of the action of $\Gamma_{e}$ on $\operatorname{Ind}_{F}(H)$ appears in $e\cdot V$. Indeed, suppose that $W_{1},W_{2}\in\operatorname{Ind}_{L}(H)$ both appear in $e\cdot V$, so $e\cdot V=W_{1}\oplus W_{2}\oplus U$ for some $U\in\operatorname{\mathbf{Rep}}_{L}(H)$. Then $(L\rtimes G)\otimes_{L[H]}(e\cdot V)\cong V^{\oplus|\Gamma_{e}|}$ by Corollary 8.5, and so by Lemma 2.14 both $(L\rtimes G)\otimes_{L[H]}W_{1}$ and $(L\rtimes G)\otimes_{L[H]}W_{2}$ are isomorphic to a direct sum of a number of copies of $V$. In particular, there are $a,b\geq 1$ with
$\left((L\rtimes G)\otimes_{L[H]}W_{1}\right)^{\oplus a}\cong\left((L\rtimes G)\otimes_{L[H]}W_{2}\right)^{\oplus b}$
@@
$\bigoplus_{x\in\Gamma_{e}}(x\cdot W_{1})^{\oplus a}\cong\bigoplus_{y\in\Gamma_{e}}(y\cdot W_{2})^{\oplus b}$
by Proposition 8.17 and therefore $W_{2}\cong\gamma\cdot W_{1}$ for some $\gamma\in\Gamma_{e}$, by Lemma 2.14.
For the remaining claims, suppose that $W\in\operatorname{Ind}_{F}(H)$. From Proposition 8.17, $W$ is a direct summand of $e\cdot((L\rtimes G)\otimes_{L[H]}W)$, and thus of $e\cdot V$ for some $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ after writing $(L\rtimes G)\otimes_{L[H]}W$ as a direct sum of indecomposable objects and using Lemma 2.14. For this $V$, from the above we have that
$V^{|\Gamma_{e}|}\cong(L\rtimes G)\otimes_{L[H]}(e\cdot V)\cong\bigoplus_{\gamma\in\Gamma_{e}/\Gamma_{e,W}}((L\rtimes G)\otimes_{L[H]}(\gamma*W))^{m(V)}$
and therefore $(L\rtimes G)\otimes_{L[H]}(\gamma*W)=V^{\oplus r}$ for any $\gamma\in\Gamma_{e}$ by Lemma 2.14, where $r=|\Gamma_{e,W}|/m(V)$.
In particular, the $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ for which $W$ is a direct summand of $e\cdot V$ is unique. Finally, from Corollary 8.14 we have that $V$ is irreducible if and only if $W$ is irreducible. ∎
We can use this to characterise when $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple, improving on Proposition 2.11. It also possible to show this directly, using a generalisation of Maschke’s Theorem [KUN, Lem. 1.3].
Corollary 8.21.
Suppose that $G$ is finite. Then $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple if and only if $|H|\in L^{\times}$.
Proof.
By Lemma 2.14, the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple if and only if $\operatorname{Irr}_{L}^{\rtimes}(G)=\operatorname{Ind}_{L}^{\rtimes}(G)$, which by Theorem 8.20 is equivalent to the statement that $\operatorname{Irr}_{F}(H)=\operatorname{Ind}_{F}(H)$, or in other words that $\operatorname{\mathbf{Rep}}_{F}(H)$ is semisimple. By Maschke’s Theorem this is equivalent to $|H|$ being invertible in $F$, which is the same as $|H|$ being invertible in $L$, as $L$ is the product of copies of $F$. ∎
In general, we will see that any positive integer can occur as some $m(V)$ (Section 16). One case when we know that the numbers $m(V)$ are equal to $1$ is the following. This applies in particular when $K$ is a finite field, by Wedderburn’s Little Theorem [milneCFT, Thm. IV.4.1].
@@
and thus $\operatorname{End}_{F[H]}(e\cdot V)$ is commutative because $\operatorname{End}_{L\rtimes G}(V)$ is a field. In particular, $e\cdot V$ cannot contain any factors with multiplicity greater than $1$, so $m(V)=1$. ∎
@@ 8.6. The Case When $L$ is a Field
$\begin{array}[]{cc}\begin{aligned} \operatorname{Ind}_{L}^{\rtimes}(G)&\xleftrightarrow{\sim}\operatorname{Ind}_{L}(H)/\Gamma,\\ \operatorname{Ind}_{L}^{\rtimes}(S)&\xleftrightarrow{\sim}\operatorname{Ind}_{L}(H)/\Gamma_{S}.\end{aligned}&\begin{aligned} \operatorname{Irr}_{L}^{\rtimes}(G)&\xleftrightarrow{\sim}\operatorname{Irr}_{L}(H)/\Gamma,\\ \operatorname{Irr}_{L}^{\rtimes}(S)&\xleftrightarrow{\sim}\operatorname{Irr}_{L}(H)/\Gamma_{S}.\end{aligned}\end{array}$
Theorem 8.20 further provides a description of induction and restriction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$, and induction and restriction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$. We now wish to understand induction and restriction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S)$.
@@
Let $\mathcal{A}$} be an orbit of $G$ on $\operatorname{Ind}_{L}(H)$, corresponding to $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$, and let
@@
be a decomposition of $\mathcal{A}$} into $S$-orbits. Let $U_{i}\in\operatorname{Ind}_{L}^{\rtimes}(S)$ correspond to ${\mathcal{A}}_{i}$, so
@@
for some $m(U_{i})\geq 1$ (by Theorem 8.20). Then $m(U_{i})\mid m(V)$, $m(V)\mid m(U_{i})\cdot[\Gamma_{X}:\Gamma_{S,X}]$,
$V|_{S}\cong\bigoplus_{i=1}^{r}U_{i}^{\oplus\frac{m(V)}{m(U_{i})}},$
@@
for any $X\in{\mathcal{A}}_{i}$. Furthermore, $V$ is irreducible if and only if each $U_{i}$ is irreducible.
@@
For the first claim, note that using Corollary 8.6 for $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S)$,
$(V|_{S})^{m(U_{1})\cdots m(U_{r})}\cong\bigoplus_{i=1}^{r}U_{i}^{\oplus m(V)m(U_{1})\cdots m(U_{i-1})m(U_{i+1})\cdots m(U_{r})}$
because their restrictions to $H$ are isomorphic. Then the description of $V|_{S}$ follows from Lemma 2.14.
For the second claim, take any $X\in{\mathcal{A}}_{i}$, and note that by Theorem 8.20
$V^{\oplus\frac{|\Gamma_{X}|}{m(V)}}\cong\operatorname{Ind}_{H}^{G}(X)\cong\operatorname{Ind}_{S}^{G}(\operatorname{Ind}_{H}^{S}(X))\cong\operatorname{Ind}_{S}^{G}(U_{i})^{\oplus\frac{|\Gamma_{S,X}|}{m(U_{i})}}.$
The description of $\operatorname{Ind}_{S}^{G}(U_{i})$ then simiarly follows directly from Lemma 2.14.
For the final claim, note that $V$ is irreducible if and only if each element of $\mathcal{A}$} is irreducible, and $U_{i}$ is irreducible if and only if each element of ${\mathcal{A}}_{i}$ is irreducible, in each case by Theorem 8.20. ∎
@@
The main case of interest is the following. To state it, note that if $S$ is a normal subgroup of $G$, then there is an action of $G/S$ on $\operatorname{Ind}_{L}^{\rtimes}(S)$, where for $g\in G$, and $U\in\operatorname{Ind}_{L}^{\rtimes}(S)$, gUg*U is the abelian group $U$ but with $L\rtimes S$-module structure
@@
for $\lambda\in L$, $s\in S$ and $u\in U$. This action preserves the subset $\operatorname{Irr}_{L}^{\rtimes}(S)$ of $\operatorname{Ind}_{L}^{\rtimes}(S)$.
@@
The action of $G/H$ on $\operatorname{Ind}_{L}(H)$ is the restriction to the diagonal of an action of the product $G/H\times G/H$ on $\operatorname{Ind}_{L}(H)$ (where one copy acts by conjugation in $G$ and the other acts on scalars - see Section 14). When $S>H$, however, this action of $G/S$ on $\operatorname{Ind}_{L}^{\rtimes}(S)$ is not the restriction of an action of the product $G/S\times G/S$ in the same way: one only has an action of the diagonal.
@@
Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$, and that for some corresponding $W\in\operatorname{Ind}_{L}(H)$, $G_{W}$ is normal in $G$. Set $G_{V}\coloneqq G_{W}$, which is independent of the choice of $W$. Let $U\in\operatorname{Ind}_{L}^{\rtimes}(G_{V})$ correspond to $W$. Then
@@
$V$ $\cong\operatorname{Ind}_{G_{V}}^{G}(g*U),$
$V$ is irreducible if and only if $U$ is irreducible, and for any $g\in G$, gUg*U corresponds to gWg*W, so
@@
Now we are interested in the condition that $L$ is furthermore a splitting field for $W\in\operatorname{Ind}_{L}(H)$ (meaning that $\operatorname{End}_{L[H]}(W)=L$, or equivalently that $W_{L^{\prime}}$ is irreducible for any field extension $L^{\prime}/L$).
@@
Suppose that $W\in\operatorname{Ind}_{L}(H)$, corresponding to $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$. Then the $K$-algebra $\operatorname{End}_{L\rtimes G}(V)$ is central over $K$ if and only if $\Gamma$ fixes $W$ and $Z(\operatorname{End}_{L[H]}(W))=L$.